OpenStudy (anonymous):

The second derivative of (4-x^2)*exp(-1.16x^2)

OpenStudy (anonymous):

can anyone help me with this?? =]

OpenStudy (anonymous):

you can use the product rule right?

OpenStudy (anonymous):

h(x)=f(x)g(x) h(x)=f(x)g(x)=f(x)g(x)+f(x)g(x)

OpenStudy (anonymous):

h'(x) = [f(x)g(x)]' = f'(x)g(x)+f(x)g'(x)

OpenStudy (anonymous):

yes but it doesn't work. my online program does not accept it. (4-x^2)(5.3824*^2*exp(-1.16x^2)-(-2.32xexp(-1.16x^2)(-2x)-(exp(-1.16x^2)(-2)-(-2x)(-2.32xexp(-1.16x^2)

OpenStudy (anonymous):

is that what you got?

OpenStudy (anonymous):

I didn't work it out. I was just pointing out that you can use the product rule for this kind of problem.

OpenStudy (anonymous):

i did. do you think you can check my work?

OpenStudy (anonymous):

but your solution seems correct. see if you can simplify it further. Maybe your online tool will accept it then

OpenStudy (anonymous):

I think I am getting a slightly different answer from you. check your calculations.

OpenStudy (anonymous):

what is slightly different about it? I just double checked!

OpenStudy (anonymous):

did you simplify it. wait let me post my solution

OpenStudy (anonymous):

you can set x^2 = y and factor the second term.

OpenStudy (anonymous):

as a quadratic equation of course.